OpenStudy (anonymous):

Part A: Solve C = (F - 32) for F. (4 points) Part B: Determine the value of F when C = 8 °C. (2 points) Part C: Solve -np - 80 > 60 for n. Show your work.

OpenStudy (gabebae):

Part A: C= (F-32) +32 +32 --------- C+32= F Part B: 8+32= F 40 = F Part C: -np - 80 > 60 +80 +80 ------------- -np > 140 -- -- p p ------------- -n > 140/p or 140 over p just so you don't get confused